# Neutral-current background induced by atmospheric neutrinos at large liquid-scintillator detectors: III. Quantitative calculations for reactor neutrinos

### v1

@@ 4 NC Background Predictions at Low Visible Energies
| NC background rate without deexcitation (ktyr) | |||
@@
| NC background rate with deexcitation (ktyr) | |||
@@
Fig. 10 presents event rates with different final-state nuclei across distinct prompt energy ranges of (0.7 100) MeV, (0.7 12) MeV, and (12 30) MeV, where the last two energy regions are associated with the detection regions of reactor neutrino and DSNB signals, respectively. The predominance of the NC background with C is evident. The deexcitation effect minimally impacts the overall event rate, causing a reduction in rates for dominant channels with C and B, while increasing rates for channels with lighter nuclei such as Be, Li and Be. When comparing the new and old neutrino interaction models, the rates of the new models do not surpass those of the old models, with only slightly higher rates observed for new GENIE’ models in the (0.7 12) MeV energy range.
@@ 4.1 Reactor
Reactor originates from the -decays of neutrino-rich fission fragments, primarily arising from four fission isotopes: U, U, Pu, and Pu. Their detection takes place through the IBD reaction. Taking the JUNO detector as an example, we have calculated the IBD rate and spectrum within the energy range extending from 0.7 MeV to 12 MeV [7]. The expected IBD signal rate, accounting for the oscillation effect, is 1514.8 yrkt.
The NC background spectra and rates for both exclusive channels and the inclusive one can be extracted from Fig. 9 and Fig. 10. From Tab. 2, the NC background rates exhibit a notable similarity between old and new models. This similarity arises from the fact that the primary distinction between old and new models lies in the setting of 1.35 GeV in Model-N3 for the old models, which mainly affects the higher energy range, as mentioned earlier. Additionally, the NC background rates experience a reduction of around 13% for both old and new models in this visible energy range due to the deexcitation effect. The associated uncertainty, represented by the model variations, ranges from around 26% to 30% when comparing rates without deexcitation to rates with deexcitation. The dominant source of uncertainty is attributed to the variation in nuclear effect models. For conservative purposes, the maximum NC background level is derived from the GENIE model, specifically 2.9 kt yr from Model-G1 in old models and 3.1 kt yr from Model-G4 in new models.
@@ 4.2 DSNB
Based on our discussions in prior papers [1, 7], we have determined the optimized observed prompt energy range for DSNB discovery potential to be from 12 MeV to 30 MeV. In this energy range, the typical predicted rate of DSNB IBD signals in LS detectors is around (0.1 0.2) kt yr. The prediction of the NC background within this energy range has been performed using the old neutrino interaction models. By averaging six old model calculations for prediction and considering model variations as uncertainty, we obtained a value of (3.0 0.5) kt yr for the NC background in the prompt energy range from 12 to 30 MeV. In this work, we have updated the prediction with new neutrino generator models. The average rates with model variations for both old and new models are shown in Tab. 2. Since we do not use a higher value of in the new models, the average rates becomes lower, and the model variation becomes smaller in this energy range. Additionally, the associated uncertainty, represented by the model variations, ranges from around 12% to 14% when comparing rates without deexcitation to rates with deexcitation.
@@ 5 Summary
In this paper, we have employed the methodology outlined in the previous paper [1] for systematically calculating the NC background induced by atmospheric neutrino interactions with C nuclei in the LS detectors. As our understanding of neutrino-nucleus interaction models and neutrino generators has advanced, we have incorporated five contemporary neutrino generator models from GENIE (3.0.6) and NuWro (19.02) for the calculations. In the present work, our focus is on processes relevant to the experimental searches for the IBD signals of in the visible energy range below 100 MeV, specifically for reactor and DSNB. In this energy range, the QE and processes of neutrino-C interactions are known to be the most important. We consider model variations related to the -nucleon NC interactions and final-state information of QE and , and investigate the impact of the nuclear model, FSI, nuclei deexcitation, and secondary interactions on the final-state information.
For the detection of reactor with visible energy lower than 12 MeV, utilizing all new neutrino generator models from two neutrino event generators (GENIE and NuWro) along with TALYS for deexcitation, we have determined that the event rate of the IBD-like NC background is (2.1 0.7) kt yr. This rate is comparable to the result obtained using precedent neutrino generator models. The dominant source of uncertainty is attributed to variations in nuclear effect models, and the deexcitation process increases the uncertainty from 26% to 30%. The NC background contributes to around 0.2% of the reactor signal, remaining one of the non-negligible backgrounds for precise measurements.
For the detection of DSNB in LS detectors, the focus is on the narrow window of visible energy in the range of (12 30) MeV. We have updated the NC background prediction by incorporating all new neutrino generator models. The event rate is (3.0 0.5) kt yr with old neutrino generator models, which is updated to (2.9 0.4) kt yr with new neutrino generator models. Further reduction of the NC background level and uncertainty can be achieved by implementing PSD and employing measurements in LS detectors.
Our calculation in this work will help estimate single, double-coincident and triple-coincident backgrounds. Further measurements from JUNO can refine model calculations, including nuclear models, FSIs, and secondary interactions in LS. This comprehensive approach will enhance our understanding of neutrino interactions and contribute to refining models for various experimental settings [17, 52, 53, 54].
@@ Acknowledgements
The authours would like to thank Wan-Lei Guo and Xian-Guo Lu for helpful discussions. We also thank Davide Basilico and Liang Zhan for carefully reading the manuscript and useful comments. This work is supported in part by National Natural Science Foundation of China under Grant Nos. 12075255, 11835013, 12125506.

### v2

@@ 4 NC Background Predictions at Low Visible Energies
| NC background rate without deexcitation (ktyr)-1 | |||
@@
| NC background rate with deexcitation (ktyr)-1 | |||
@@
Fig. 10 presents event rates with different final-state nuclei across distinct prompt energy ranges of (0.7 100) MeV, (0.7 12) MeV, and (12 30) MeV, where the last two energy regions are associated with the detection regions of reactor neutrino and DSNB signals, respectively. The predominance of the NC background with 11C is evident. The deexcitation effect minimally impacts the overall event rate, causing a reduction in rates for dominant channels with 11C and 10B, while increasing rates for channels with lighter nuclei such as 9Be, 6Li and 8Be. When comparing the new and old neutrino interaction models, the rates of the new models do not surpass those of the old models, with only slightly higher rates observed for new GENIE’ models in the (0.7 12) MeV energy range.
@@ 4.1 Reactor
Reactor originates from the -decays of neutrino-rich fission fragments, primarily arising from four fission isotopes: 235U, 238U, 239Pu, and 241Pu. Their detection takes place through the IBD reaction. Taking the JUNO detector as an example, we have calculated the IBD rate and spectrum within the energy range extending from 0.7 MeV to 12 MeV [26]. The expected IBD signal rate, accounting for the oscillation effect, is 1514.8 yr-1kt-1.
The NC background spectra and rates for both exclusive channels and the inclusive one can be extracted from Fig. 9 and Fig. 10. From Tab. 2, the NC background rates exhibit a notable similarity between old and new models. This similarity arises from the fact that the primary distinction between old and new models lies in the setting of 1.35 GeV in Model-N3 for the old models, which mainly affects the higher energy range, as mentioned earlier. Additionally, the NC background rates experience a reduction of around 13% for both old and new models in this visible energy range due to the deexcitation effect. The associated uncertainty, represented by the model variations, ranges from around 26% to 30% when comparing rates without deexcitation to rates with deexcitation. The dominant source of uncertainty is attributed to the variation in nuclear effect models. For conservative purposes, the maximum NC background level is derived from the GENIE model, specifically 2.9 kt-1 yr-1 from Model-G1 in old models and 3.1 kt-1 yr-1 from Model-G4 in new models.
@@ 4.2 DSNB
Based on our discussions in prior papers [1, 26], we have determined the optimized observed prompt energy range for DSNB discovery potential to be from 12 MeV to 30 MeV. In this energy range, the typical predicted rate of DSNB IBD signals in LS detectors is around (0.1 0.2) kt-1 yr-1. The prediction of the NC background within this energy range has been performed using the old neutrino interaction models. By averaging six old model calculations for prediction and considering model variations as uncertainty, we obtained a value of (3.0 0.5) kt-1 yr-1 for the NC background in the prompt energy range from 12 to 30 MeV. In this work, we have updated the prediction with new neutrino generator models. The average rates with model variations for both old and new models are shown in Tab. 2. Since we do not use a higher value of in the new models, the average rates become lower, and the model variation becomes smaller in this energy range. Additionally, the associated uncertainty, represented by the model variations, ranges from around 12% to 14% when comparing rates without deexcitation to rates with deexcitation.
@@ 5 Summary
In this work, we have extended the methodology of predicting the NC interactions by boosting its completeness and applicability in the neutrino interactions and the propagation of final-state particles in large LS detectors. Several outstanding advancements are achieved including the incorporation of five new data-driven models from GENIE (3.0.6) and NuWro (19.02) and the secondary interactions in the LS medium. With a focus on deepening our understanding of GeV neutrino interactions, our calculations keep pace with the persistent optimization and validation of neutrino interaction models. Moreover, a GEANT4-based detector simulation has been formally integrated into our calculations, paving the way for the examination of secondary interactions of final-state particles. The NC interactions are significant backgrounds for the experimental searches of the IBD signals of in the visible energy less than 100 MeV, specifically for reactor and geo , and the DSNB. Within this energy range, the QE and processes of neutrino-12C interactions are recognized as the paramount. We have examined the model variation associated with the neutrino nucleon initial interactions and final-state information in QE and processes, and shown that the impact of nuclear ground state, FSI, nuclear deexcitation on inclusive and exclusive cross-sections, and the effect of secondary interactions on neutron multiplicity are ascertained.
This work provides the first detailed prediction for the NC backgrounds in the energy range below 100 MeV. The event rate of the IBD-like NC background with a visible energy below 12 MeV is estimated as (2.1 0.7) kt-1 yr-1. The dominant source of uncertainty is attributed to variations in models of the nuclear effect. The NC background accounts for around 0.2% for the reactor IBD signal at JUNO [25], remaining as one of the non-negligible backgrounds for precise measurements. This prediction has been utilized in determining the neutrino mass ordering with reactor . Additionally, this prediction proves to be useful in searches for geo-neutrinos. Regarding the substantial interest in the first detection of the DSNB in the energy range between 12 and 30 MeV, we have updated the prediction of the corresponding NC background by using advanced data-driven models, in which the event rate is (2.9 0.4) kt-1 yr-1.
Our methodology for predicting the NC interactions provides a wide array of application scenarios. Firstly, it is instrumental not only in estimating the double-coincident NC background, as studied in this work, but also in calculating single, triple-coincident, or other types of NC induced backgrounds. Secondly, it can be easily adapted for predicting charged-current interactions. Thirdly, while currently applied to LS detectors, this method is also suitable for a variety of experimental settings, such as Hyper-K [59], Theia [60], and DUNE [61], for forecasting their related observations. Finally, we would like to note that future measurements from these experiments could potentially test and refine the current model predictions, including aspects of primary neutrino-nucleus scattering and secondary interactions in the detectors. We plan to explore this intriguing study in future publications.
@@ Acknowledgements
The authors would like to thank Wan-Lei Guo and Xian-Guo Lu for helpful discussions. We also thank Davide Basilico and Liang Zhan for carefully reading the manuscript and useful comments. This work is supported in part by National Natural Science Foundation of China under Grant Nos. 12075255, 11835013, 12125506.